HOW TO SOLVE------------------------------------5000(1 + 10 * 3/100) ......................PLEASE HELP >:/
1
Expert's answer
2013-01-10T07:20:13-0500
You have to keep in mind priority of mathematical operations when solving problems like this. Thus, if you have expression like one in braces here, first multiplication/division is performed and then addition:
5000(1 + 10 * 3/100)=5000(1 + 3/10)=5000*1.3=6500